What is the probability that a random sample of 70 pregnancies has a mean gestation period of 124 days or​ less?
1
Expert's answer
2020-05-18T19:16:14-0400

Suppose the lengths of the pregnancies of a certain animal are approximately normally distributed with mean "\mu=120" days and standard deviation "\sigma=10" days. 

What is the probability that a random sample of 70 pregnancies has a mean gestation period of 124 days or​ less?

Let "\bar{X}=" a mean gestation period: "\bar{X}\sim N(\mu, \sigma^2/n)." Then


"Z={\bar{X}-\mu\over \sigma/\sqrt{n}}\sim N(0, 1)"

Given that "\mu=120, \sigma=10, n=70."


"P(\bar{X}\leq124)=P(Z<{124-120\over 10/\sqrt{70}})\approx"

"\approx P(Z\leq3.346640)\approx0.999591"